OpenStudy (anonymous):

Susan rented a flat at $200 per week. She also had to pay a deposit of $1400. Part A: Write a function to show the total amount of money f(x) that Susan has to pay for renting the flat for x weeks. (4 points) Part B: What is the total rent that Susan has to pay for using the flat for 12 weeks? (4 points) Part C: If the deposit amount decreases by $100, what would be the total rent for 6 weeks? (2 points) Part A: F(x) = Part B: What is the total rent that Susan has to pay for using the flat for 12 weeks? Plug and chug! Part C: If the deposit amount decreases by $100, w

OpenStudy (anonymous):

@perl @aum

OpenStudy (anonymous):

Do you still need the answer?

OpenStudy (anonymous):

yes please @blw0

OpenStudy (anonymous):

@blw0

OpenStudy (anonymous):

@blw0

OpenStudy (anonymous):

Sorry I was helping someone else I will start.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Part A: f(x) = 200x + 1400

OpenStudy (anonymous):

Part B: f(x) = 200(12) + 1400 = 3800

OpenStudy (anonymous):

Part C: f(x) = 200(6) + 1300 = 2500

OpenStudy (anonymous):

thats all?

OpenStudy (anonymous):

Yeah do you have multiple choice?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but thank you that was a lot easier than i had thought

OpenStudy (anonymous):

The function is f(x) = 200(w) + b where w is the amount of weeks and b is the deposit. It is all self explanatory past that